### v1

# Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning

Abstract
Low-Rank Adaptation (LoRA) has become ubiquitous for efficiently fine-tuning foundation models. However, federated fine-tuning using LoRA is challenging due to suboptimal updates arising from traditional federated averaging of individual adapters. Existing solutions either incur prohibitively high communication cost that scales linearly with the number of clients or suffer from performance degradation due to limited expressivity. We introduce Federated Silver Bullet (Fed-SB), a novel approach for federated fine-tuning of LLMs using LoRA-SB, a recently proposed low-rank adaptation method. LoRA-SB optimally aligns the optimization trajectory with the ideal low-rank full fine-tuning projection by learning a small square matrix ($R$) between adapters $B$ and $A$, keeping other components fixed. Direct averaging of $R$ guarantees exact updates, substantially reducing communication cost, which remains independent of the number of clients, and enables scalability. Fed-SB achieves state-of-the-art performance across commonsense reasoning, arithmetic reasoning, and language inference tasks while reducing communication costs by up to 230x. In private settings, Fed-SB further improves performance by (1) reducing trainable parameters, thereby lowering the noise required for differential privacy and (2) avoiding noise amplification introduced by other methods. Overall, Fed-SB establishes a new Pareto frontier in the tradeoff between communication and performance, offering an efficient and scalable solution for both private and non-private federated fine-tuning. Our code is publicly available at https://github.com/CERT-Lab/fed-sb.
Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning
Raghav Singhal*1, Kaustubh Ponkshe*1, Rohit Vartak2, Lav R. Varshney3, Praneeth Vepakomma1,4 1Mohamed bin Zayed University of Artificial Intelligence, 2Duke University, 3University of Illinois Urbana-Champaign, 4Massachusetts Institute of Technology *Equal contribution. Order decided by coin toss.
1 Introduction
Large language models (LLMs) have demonstrated remarkable generalization across a wide range of tasks (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; Raffel et al., 2020). Fine-tuning (FT) remains the most effective approach for aligning LLMs to specific data distributions and reinforcing desired properties. However, as model sizes scale, full FT becomes increasingly prohibitive due to its substantial computational cost. To address this, parameter-efficient fine-tuning (PEFT) techniques, such as low-rank adaptation (LoRA, Hu et al. (2021)), have emerged as viable alternatives, offering a favorable trade-off between computational efficiency and performance. Variants of LoRA, including QLoRA (Dettmers et al., 2024), DoRA (Liu et al., 2024), AdaLoRA (Zhang et al., 2023), and LoRA-SB (Ponkshe et al., 2024), further refine this paradigm by optimizing memory efficiency, training dynamics, and generalization.
Federated learning (FL) is a popular method for training models in settings where data is siloed across multiple entities (Konečný et al., 2017; Kairouz et al., 2021; Bonawitz et al., 2019). Federated FT extends this paradigm by enabling large models, pre-trained on public data, to be efficiently adapted to private, distributed datasets without requiring clients to share their local data. Existing methods predominantly rely on LoRA-based techniques to learn client-specific adaptations (Zhang et al., 2024b). However, optimizing federated aggregation often involves tradeoffs between model performance (Sun et al., 2024) and communication efficiency (Wang et al., 2024; Singhal et al., 2025), necessitating careful design choices to balance these competing objectives.
Figure 1: Performance vs. communicated parameter cost (log scale) for Fed-SB and other federated fine-tuning methods in both non-private and privacy-preserving federated settings. Fed-SB advances the performance-communication cost Pareto frontier across all models and tasks, achieving state-of-the-art accuracy while significantly reducing communication cost. Communicated parameters are in thousands for BERT and millions for other models.
LoRA-SB (Ponkshe et al., 2024), a state-of-the-art approach, optimally simulates full fine-tuning in low-rank spaces by learning an $r\times r$ matrix between the low-rank adapters $\mathbf{A}$ and $\mathbf{B}$ while keeping other components fixed. This design reduces trainable parameters and enables better updates through its initialization strategy. Moreover, LoRA-SB demonstrates that this optimal approximation is not achievable with standard LoRA-based methods. LoRA-SB learns higher-rank updates with 2–4x greater rank than LoRA while requiring 45-90x fewer parameters. We propose Fed-SB, a federated variant of LoRA-SB, providing an ideal framework for (private) federated FT. Fed-SB overcomes limitations in LoRA-based federated FT while being significantly more computation- and communication-efficient. Notably, it enables exact and optimal aggregation by simply averaging the learnable matrix $\mathbf{R}$.
Differential privacy (DP) is a well-established framework for ensuring strong privacy guarantees (Dwork, 2006; Dwork et al., 2014), which is particularly crucial in federated settings. DP-SGD is a widely used privacy-preserving optimization method (Abadi et al., 2016), but its challenges are exacerbated in federated FT, where noise injected for privacy amplifies divergence across client models (Sun et al., 2024). Learning in DP-SGD is more effective when the number of learnable parameters is reduced, as the magnitude of noise added for privacy guarantees scales with the parameter count. Fed-SB mitigates this issue to yield improved performance, since it inherently has fewer learnable parameters and thus less noise injection. Furthermore, we show that Fed-SB avoids noise amplification introduced by other methods, further enhancing privacy-preserving learning.
Fed-SB pushes the performance vs communication cost Pareto frontier, offering an extremely efficient and scalable solution for both private and non-private federated FT, as shown in Figure 1. It consistently has superior performance while substantially reducing communication overhead than other methods. Our key contributions are summarized as follows.
  • We propose Fed-SB, a federated fine-tuning method that achieves exact and optimal aggregation in low-rank adaptation without incurring prohibitive communication costs or performance degradation.
  • Fed-SB significantly reduces communication cost—by up to 230x compared to state-of-the-art methods—by requiring only an $r\times r$ matrix to be transmitted per aggregation.
  • We demonstrate that Fed-SB is particularly well-suited for privacy-preserving federated fine-tuning, as it minimizes noise by reducing the number of learnable parameters and leveraging linearity in the aggregate update.
  • Extensive experiments on four models across three diverse benchmarks show that Fed-SB consistently outperforms existing methods while drastically reducing communication overhead in both private and non-private federated settings, establishing a new Pareto frontier in federated fine-tuning.
FedIT FLoRA FedEx-LoRA FFA-LoRA Fed-SB
Exact aggregation
Learnable params. / client $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}((m+n)r)$ $\mathcal{O}(mr)$ $\mathcal{O}(r^{2})$
Communication cost $\mathcal{O}((m+n)r)$ $\mathcal{O}(min(c(m+n)r,mn))$ $\mathcal{O}(min(c(m+n)r,mn))$ $\mathcal{O}(mr)$ $\mathcal{O}(r^{2})$
No noise amplification
Privacy benefit (less params.)
Optimal expressivity
Table 1: Advantages of Fed-SB over various state-of-the-art federated fine-tuning methods ($c$ clients).
2 Preliminaries and Motivation
Federated Fine-Tuning. Given a pretrained weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the objective in FT is to learn an update $\Delta\mathbf{W}$ for a given dataset. LoRA (Hu et al., 2021) remains the preferred method, where low-rank adapter matrices $\mathbf{A}\in\mathbb{R}^{r\times n}$ and $\mathbf{B}\in\mathbb{R}^{m\times r}$ are learned such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$. In federated learning, the dataset is distributed across $c$ clients, and the goal is to learn $\Delta\mathbf{W}$ without sharing local data with a central server. To achieve this, each client learns its own adapter matrices $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$. The server aggregates these updates to refine $\mathbf{W}$, along with globally beneficial representations of $\mathbf{A}$ and $\mathbf{B}$, ultimately producing a shared aggregate model $\mathbf{W}^{\text{agg}}$. Next, each client continues the local FT process, followed by aggregation at the end of each round. This cycle repeats over multiple rounds. We summarize some of the state-of-the-art federated FT methods below.
Fed-IT (Zhang et al., 2024b) updates the adapters $\mathbf{A}$ and $\mathbf{B}$ using the standard FedAvg (McMahan et al., 2017) algorithm:
$\mathbf{A}^{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i},\quad\mathbf{B}^{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}.$ (1)
FedEx-LoRA (Singhal et al., 2025) follows the same aggregation but introduces an additional error correction matrix $\mathbf{W}_{\text{err}}$ of rank $\min(cr,m,n)$:
$\mathbf{W}_{\text{err}}=(\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i}\mathbf{B}_{i})-(\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i})(\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}).$ (2)
FLoRA (Wang et al., 2024) follows the same principle as FedEx-LoRA but achieves it by stacking the adapter matrices, and reinitializes them randomly at the end of each communication round.
FFA-LoRA (Sun et al., 2024) keeps $\mathbf{A}$ fixed while training (and aggregating) only $\mathbf{B}$ matrices.
$\mathbf{B}^{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}.$ (3)
(Approximate) Differential Privacy. DP, introduced by Dwork (2006), is a widely adopted mathematical framework for privacy preservation. A randomized mechanism $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$, mapping a domain $\mathcal{D}$ to a range $\mathcal{R}$, satisfies $(\epsilon,\delta)$-differential privacy if, for any two adjacent inputs $d,d^{\prime}\in\mathcal{D}$ and any subset of outputs $S\subseteq\mathcal{R}$, the following holds:
$\operatorname{Pr}[\mathcal{M}(d)\in S]\leq e^{\epsilon}\operatorname{Pr}[\mathcal{M}(d^{\prime})\in S]+\delta.$ (4)
DP-SGD. DP-SGD (Abadi et al., 2016) is a privacy-preserving variant of stochastic gradient descent (SGD) designed to ensure DP during model training. It enforces privacy by clipping per-sample gradients to a fixed norm $C$ to limit their sensitivity and then adding isotropic Gaussian noise $\mathcal{N}\left(0,\sigma^{2}C^{2}\mathbf{I}\right)$, where $\sigma$ controls the noise magnitude. The cumulative privacy loss over iterations is quantified using the moments accountant (Wang et al., 2019) and Rényi DP (Mironov, 2017), which offer a tight bound on the final privacy parameter $\epsilon$.
Exact Aggregation in Fed. LoRA: Tradeoff b/w Performance and Communication Costs.
Standard federated averaging of individual LoRA adapters (FedIT, Zhang et al. (2024b)) introduces inexactness in aggregation, as the ideal update should be the average of client updates.
$\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}_{i}\times\frac{1}{c}\sum_{i=1}^{c}\mathbf{A}_{i}}_{\text{Vanilla aggregation in LoRA (FedIT)}}\\ \neq\underbrace{\mathbf{W}_{0}+\frac{1}{c}\sum_{i=1}^{c}(\mathbf{B}_{i}\mathbf{A}_{i})}_{\text{Ideal aggregation}}.$ (5)
The inexactness arises because the ideal averaged updates, given by $\sum_{i=1}^{c}\mathbf{B}_{i}\mathbf{A}_{i}$, often exceed rank $r$, violating the low-rank constraint imposed by LoRA. To address this, FedEx-LoRA and FLoRA introduce $\mathbf{W}_{\text{err}}$ as a higher-rank correction term within the pre-trained weight matrix $\mathbf{W}_{0}$, which is inherently high-rank. This correction ensures exact aggregation, leading to consistently improved performance over FedIT.
This, however, comes at the cost of increased communication. Since the error matrix is high rank, it substantially increases the amount of data transmitted per round. The communication cost is determined by the number of parameters sent during aggregation, which, for an $m\times n$ matrix, is proportional to its rank. As a result, in FedEx-LoRA and similar methods that enforce exact aggregation, communication cost scales linearly with the number of clients relative to Fed-IT. This becomes particularly concerning when the number of clients grows large, potentially requiring the transmission of the entire model’s weights.
FFA-LoRA addresses inexact aggregation by keeping only $\mathbf{B}$ trainable while fixing $\mathbf{A}$ uniformly across clients. However, this comes at the cost of reduced expressivity and limits the benefits of jointly optimizing $\mathbf{A}$ and $\mathbf{B}$. As a result, performance degrades, as demonstrated previously (Singhal et al., 2025). This stems from two factors: suboptimal individual updates and the need for higher-rank adaptations. Freezing $\mathbf{A}$ leads to suboptimal updates, even in centralized training, where FFA-LoRA underperforms compared to LoRA. Additionally, Mahla et al. (2025) show that models trained using FFA-LoRA progressively deviate from the optimal hypothesis. Empirical evidence shows that the advantages of exactness are outweighed by the degradation caused by these factors.
Private Fine-Tuning. Pre-training on public data followed by FT on user-specific private data [^1] Although pre-training data may be publicly available, it often contains sensitive or proprietary information, raising privacy concerns. However, any privacy loss from pre-training has already occurred upon the model’s release. is a common approach for adapting models under privacy constraints (Yu et al., 2021; Tang et al., 2024). This two-stage process enhances performance in private learning while preserving user data privacy. FL naturally improves privacy by keeping data decentralized. However, even without direct data sharing, client-specific model updates can still leak sensitive information (Truong et al., 2021). Thus, developing privacy-preserving FT methods for FL is essential to ensure strong privacy guarantees while maintaining performance.
Training a model with DP-SGD introduces noise into the gradient, and consequently, into the model update itself. In the case of LoRA, this deviation from the ideal update is more pronounced than in full FT due to second-order noise terms. To illustrate this, let $\mathbf{A}$ and $\mathbf{B}$ represent the adapter updates learned without privacy. Under DP-SGD, these updates are perturbed by noise terms $\boldsymbol{\xi}_{A}$ and $\boldsymbol{\xi}_{B}$, respectively. The difference between the ideal update $\Delta\mathbf{W}$ and the noisy update $\Delta\mathbf{W}_{DP}$ is:
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\left(\mathbf{B}+\boldsymbol{\xi}_{B}\right)\left(\mathbf{A}+\boldsymbol{\xi}_{A}\right)-\mathbf{B}\mathbf{A}\\ =\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}+\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}.$ (6)
The first-order noise term, $\boldsymbol{\xi}_{B}\mathbf{A}+\mathbf{B}\boldsymbol{\xi}_{A}$, is expected and occurs even in full FT with DP-SGD. However, the second-order noise term, $\boldsymbol{\xi}_{B}\boldsymbol{\xi}_{A}$, causes noise amplification, leading to further performance degradation in LoRA-based methods (Sun et al., 2024). This issue is exacerbated in FL, as individual client updates deviate even further from the ideal global update. FFA-LoRA avoids this problem by freezing $\mathbf{A}$, preventing the introduction of additional noise terms.
A Silver Bullet Indeed. The bilinear parameterization in LoRA introduces two key challenges: inexact aggregation and noise amplification. FedEx-LoRA/FLoRA addresses the inexactness issue by enabling exact aggregation, but at the cost of a communication overhead that scales prohibitively with the number of clients. FFA-LoRA mitigates inexact aggregation and excessive communication but sacrifices performance, as it operates in a low-rank space and has reduced expressivity. An ideal method would efficiently learn higher-rank updates while inherently enabling exact aggregation without increasing communication costs. However, any LoRA-based formulation that attempts to resolve all these challenges must inevitably trade off expressivity, ultimately compromising performance. We prove that LoRA-SB provides an optimal reparameterization of the updates, effectively overcoming all limitations of LoRA in both non-private and privacy-preserving federated settings.
3 Method
Figure 2: Fed-SB: Our method achieves optimal exact aggregation by averaging only the $r\times r$ matrices $\mathbf{R}_{i}$, significantly reducing communication costs.
LoRA-SB for Fine-Tuning. LoRA-SB (Ponkshe et al., 2024) optimally approximates full FT gradients in low-rank spaces and demonstrates that its entire optimization trajectory aligns with the ideal low-rank projection of the full FT path. To achieve this, LoRA-SB fixes $\mathbf{A}$ and $\mathbf{B}$ while introducing a new trainable adapter $\mathbf{R}$ of size $r\times r$. Since $\mathbf{R}$ has rank $r$, it updates the pre-trained weight while maintaining rank $r$, making it highly parameter efficient. As a result, LoRA-SB consistently outperforms LoRA (and its variants) across various benchmarks while using 45–90x fewer trainable parameters.
Fed-SB: A Silver bullet for (Private) Federated Fine-Tuning. We propose Fed-SB, an extremely communication-efficient and high-performing federated adaptation of LoRA-SB. Instead of reparameterizing updates as a low-rank decomposition with learnable adapters, the server distributes frozen adapters $\mathbf{B}$ and $\mathbf{A}$, while clients train only a small matrix $\mathbf{R}$ (Figure 2). This enables exact aggregation, as the global update is simply the average of $\mathbf{R}$ across clients. Formally, given a pre-trained weight $\mathbf{W}_{0}$ and data distributed across $c$ clients, each client learns updates of the form:
$\Delta\mathbf{W}_{i}=\mathbf{B}\mathbf{R}_{i}\mathbf{A}.$ (7)
The server then aggregates the updates by computing the global $\mathbf{R}$ matrix:
$\mathbf{R}^{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\mathbf{R}_{i},\Delta\mathbf{W}^{\text{agg}}=\mathbf{B}\left(\frac{1}{c}\sum_{i=1}^{c}\mathbf{R}_{i}\right)\mathbf{A}.$ (8)
We show that Fed-SB effectively resolves all challenges in (private) federated FT while achieving state-of-the-art communication efficiency and performance. Table 1 highlights the advantages of Fed-SB over other methods.
Fed-SB: Exact Aggregation. Since only $\mathbf{R}$ is trainable, simple averaging of $\mathbf{R}$ across clients ensures exact aggregation without requiring updates to any other matrix. Further, the linearity of the global update with respect to the client-specific matrices $\mathbf{R}_{i}$ guarantees that exact aggregation occurs within rank $r$, preventing communication costs from scaling with the number of clients. This is because the server only needs to aggregate and transmit the $\mathbf{R}$ matrix. This can be proven directly by computing the global update $\Delta\mathbf{W}^{\text{agg}}$:
$\Delta\mathbf{W}^{\text{agg}}=\mathbf{B}\left(\frac{1}{c}\sum_{i=1}^{c}\mathbf{R}_{i}\right)\mathbf{A},$ (9)
$\quad\Delta\mathbf{W}^{\text{agg}}=\frac{1}{c}\sum_{i=1}^{c}\mathbf{B}\mathbf{R}_{i}\mathbf{A}=\frac{1}{c}\sum_{i=1}^{c}\Delta\mathbf{W}_{i}.$ (10)
Since the global update is simply the average of the individual updates, the aggregation is exact. The key advantage here is that this exact aggregation does not incur additional communication overhead like FedEx-LoRA, nor does it compromise individual update quality like FFA-LoRA.
Fed-SB: Privacy. Privacy-preserving FT with Fed-SB has two key advantages: 1) Fed-SB avoids noise amplification, which is a common issue in LoRA-based methods. 2) Since Fed-SB inherently requires fewer learnable parameters, the amount of noise added to enforce DP guarantees is significantly lower.
Avoids Noise Amplification. DP-SGD training in Fed-SB avoids second-order noise terms, as only $\mathbf{R}$ is trainable. This prevents the introduction of cross terms, thereby eliminating noise amplification. The difference between the updates with and without private training is given by:
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\left(\mathbf{R}+\boldsymbol{\xi}_{B}\right)\mathbf{A}-\mathbf{B}\mathbf{R}\mathbf{A}\\ \implies\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\boldsymbol{\xi}_{B}\mathbf{A}.$ (11)
Since the private update remains linear in $\mathbf{R}$, Fed-SB achieves the same benefits in private settings as FFA-LoRA, while avoiding its limitations.
Fewer Learnable Parameters. The noise added to gradients for DP enforcement increases with the number of trainable parameters (Bassily et al., 2014; Abadi et al., 2016; Bun et al., 2014), potentially distorting learning and degrading performance. Reducing trainable parameters improves DP performance, provided the model retains sufficient task-specific expressivity.
Lemma 1.
Consider a model with $d$ learnable parameters trained using DP-SGD. The privacy parameter $\epsilon$ for $\delta$-approximate differential privacy, given $T$ training steps and a batch size of $q$, is expressed as: $\epsilon=O(q\sqrt{Td\log(1/\delta)})=O(\sqrt{d}).$ (12)
Proof.
See Appendix A. ∎
Lemma 1 establishes that reducing the number of learnable parameters enhances privacy guarantees under the same training setup. Specifically, achieving an equivalent level of privacy requires injecting less noise per parameter when fewer parameters are trained. Since LoRA-SB optimally approximates full fine-tuning gradients, its updates remain as effective as those in LoRA while benefiting from lower noise per update, resulting in a superior privacy-utility tradeoff. More generally, any reparameterization that reduces trainable parameters leads to a smaller accumulated privacy parameter $\epsilon$, thereby improving privacy performance—provided the reduction does not compromise learning.
Fed-SB: Pushing the Pareto Frontier. Fed-SB has significantly less communication costs than other federated FT methods. This is due to two key reasons: 1) LoRA-SB achieves performance comparable to or better than LoRA while requiring 45-90x fewer trainable parameters. 2) Fed-SB aggregates only the $r\times r$ trainable matrix $\mathbf{R}$, ensuring exact aggregation without additional communication overhead. This efficiency allows Fed-SB to leverage higher-rank updates without increasing communication costs. LoRA-SB typically operates at ranks 2–4x higher than LoRA, enabling Fed-SB to capture richer model updates. Retaining high-rank information is crucial in FL (Mahla et al., 2025) and is a key factor in the superior performance of FedEx-LoRA/FLoRA over FFA-LoRA and Fed-IT beyond just aggregation exactness. We empirically validate Fed-SB’s advantages through extensive experiments.
4 Experiments & Results
Method Rank # Comm. (M) ($\downarrow$) Accuracy ($\uparrow$)
BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg
FedIT $32$ $48.63$ $62.99$ $81.50$ $73.13$ $76.83$ $71.51$ $84.89$ $70.65$ $70.62$ $74.02$
FFA-LoRA $32$ $24.31$ $62.87$ $80.03$ $68.53$ $70.02$ $65.56$ $82.95$ $66.38$ $66.85$ $70.40$
FedEx-LoRA $32$ $243.15$ $65.05$ $82.81$ $74.67$ $81.84$ $76.01$ $86.32$ $71.42$ $73.81$ $76.49$
FLoRA $32$ $243.15$ $65.05$ $82.81$ $74.67$ $81.84$ $76.01$ $86.32$ $71.42$ $73.81$ $76.49$
Fed-SB $120$ $2.83$ $64.86$ $81.66$ $74.87$ $81.67$ $75.22$ $86.03$ $70.56$ $72.25$ $75.89$
Fed-SB $160$ $5.02$ $65.57$ $82.37$ $76.15$ $84.10$ $77.98$ $86.62$ $72.10$ $73.63$ $77.32$
Fed-SB $200$ $7.85$ $\mathbf{66.66}$ $\mathbf{83.79}$ $\mathbf{77.22}$ $\mathbf{85.42}$ $\mathbf{79.56}$ $\mathbf{87.46}$ $\mathbf{72.53}$ $\mathbf{76.02}$ $\mathbf{78.58}$
Table 2: Federated fine-tuning performance of Llama-3.2 3B across eight commonsense reasoning datasets. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Model Method Rank # Comm. (M) ($\downarrow$) Accuracy ($\uparrow$)
GSM8K MATH
Mistral-7B FedIT $32$ $83.88$ $52.91$ $12.26$
FFA-LoRA $32$ $41.94$ $53.67$ $12.46$
FedEx-LoRA $32$ $2097.34$ $54.28$ $12.92$
FLoRA $32$ $2097.34$ $54.28$ $12.92$
Fed-SB $120$ $3.22$ $54.44$ $\mathbf{14.06}$
Fed-SB $160$ $5.73$ $54.81$ $13.74$
Fed-SB $200$ $8.96$ $\mathbf{56.18}$ $13.76$
Gemma-2 9B FedIT $32$ $108.04$ $74.22$ $36.30$
FFA-LoRA $32$ $54.02$ $75.06$ $35.44$
FedEx-LoRA $32$ $2701.12$ $74.68$ $36.70$
FLoRA $32$ $2701.12$ $74.68$ $36.70$
Fed-SB $120$ $4.23$ $74.75$ $36.36$
Fed-SB $160$ $7.53$ $76.88$ $36.94$
Fed-SB $200$ $11.76$ $\mathbf{77.03}$ $\mathbf{37.56}$
Table 3: Federated fine-tuning performance of Mistral-7B and Gemma-2 9B on arithmetic reasoning benchmarks GSM8K and MATH. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Overview. We evaluate across three diverse NLP benchmarks, covering models that span from BERT-base (110M) to Gemma-2 (9B), thereby encompassing both masked and autoregressive architectures. Specifically, we fine-tune Mistral-7B (Jiang et al., 2023), Gemma-2 9B (Team et al., 2024), Llama-3.2 3B (Dubey et al., 2024), and BERT-base (Devlin, 2018). Experiments are conducted on a single NVIDIA A6000 GPU (48 GB), with results averaged over three random runs. Our experiments consider both performance and communication efficiency. To optimize memory efficiency, all base models (except BERT) are loaded in torch.bfloat16. Detailed dataset specifications can be found in Appendix D. For federated data distribution, we adopt a standard protocol where client datasets are randomly sampled, following established practice in FL (Sun et al., 2024; He et al., 2020; Lai et al., 2022). Detailed settings are provided in Appendix C.
Baselines. We evaluate Fed-SB against several SOTA federated FT approaches described previously, considering both private and non-private settings. Specifically, we compare it with FedIT, FedEx-LoRA, FLoRA, and FFA-LoRA. Where applicable, we also include comparisons with standard centralized LoRA (Hu et al., 2021).
4.1 Instruction Tuning
Details. We conduct experiments in the federated non-private setting across two reasoning tasks: commonsense reasoning and arithmetic reasoning. For commonsense reasoning, we fine-tune Llama-3.2 3B on CommonSense170K—a dataset aggregating eight commonsense reasoning corpora (Hu et al., 2023)—and evaluate its effectiveness across all constituent datasets. The experiments are performed in a cross-silo federated learning setup involving $5$ clients. For arithmetic reasoning, we fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 20K samples from the MetaMathQA dataset (Yu et al., 2024) and assess their performance on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. In this setup, we distribute the federated training across $25$ clients. In both cases, we apply LoRA modules to the key, query, value, attention output, and all fully connected weights.
Results (Tables 2, 3). Our method significantly advances the performance-communication cost Pareto frontier, outperforming all previous baselines in both accuracy and communication efficiency across all models and benchmarks. Figure 3 further illustrates this significant improvement.
Method Rank # Params. (K) ($\downarrow$) Accuracy ($\uparrow$)
$\mathbf{\epsilon=1}$ $\mathbf{\epsilon=3}$ $\mathbf{\epsilon=5}$ $\mathbf{\epsilon=7.5}$ $\mathbf{\epsilon=10}$
Cent. LoRA $32$ $1181.96$ $66.49$ $67.79$ $68.17$ $70.78$ $70.81$
Cent. FFA-LoRA $32$ $592.13$ $74.40$ $75.02$ $75.02$ $76.14$ $76.60$
Cent. Fed-SB $32$ $26.88$ $73.99$ $75.09$ $74.45$ $77.01$ $76.24$
Cent. Fed-SB $48$ $57.59$ $\mathbf{75.98}$ $75.70$ $76.58$ $76.77$ $77.96$
Cent. Fed-SB $64$ $100.61$ $75.81$ $\mathbf{77.07}$ $\mathbf{77.59}$ $\mathbf{78.75}$ $\mathbf{78.08}$
Table 4: Centralized (Cent.) private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Params. denotes the number of trainable parameters. Best results are in bold.
Method Rank # Comm. (K) ($\downarrow$) Accuracy ($\uparrow$)
$\mathbf{\epsilon=1}$ $\mathbf{\epsilon=3}$ $\mathbf{\epsilon=5}$ $\mathbf{\epsilon=7.5}$ $\mathbf{\epsilon=10}$
FedIT $32$ $1181.96$ $49.57$ $51.29$ $48.53$ $55.63$ $60.96$
FFA-LoRA $32$ $592.13$ $70.11$ $71.49$ $72.69$ $73.27$ $74.02$
FedEx-LoRA $32$ $3541.26$ $67.38$ $69.68$ $72.92$ $71.89$ $74.33$
FLoRA $32$ $3541.26$ $67.38$ $69.68$ $72.92$ $71.89$ $74.33$
Fed-SB $32$ $26.88$ $70.33$ $72.68$ $73.57$ $73.62$ $73.85$
Fed-SB $48$ $57.59$ $73.7$ $74.74$ $73.66$ $74.75$ $75.02$
Fed-SB $64$ $100.61$ $\mathbf{73.83}$ $\mathbf{74.88}$ $\mathbf{76.27}$ $\mathbf{75.75}$ $\mathbf{75.86}$
Table 5: Federated private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Commonsense Reasoning (Table 2). Fed-SB ($r=200$) achieves an average accuracy improvement of 4.56% over FedIT while requiring lower communication cost. Additionally, Fed-SB ($r=200$) surpasses the previous SOTA performance methods FedEx-LoRA/FLoRA by 2.09%, while reducing communication cost by an impressive 31×. Notably, while the communication cost of FedEx-LoRA/FLoRA scales linearly with the number of clients, our method maintains a constant, client-independent communication cost. These results are obtained with just $5$ clients, implying that the full extent of our method’s communication efficiency is not fully depicted in this experiment. As the number of clients increases, the relative advantage of Fed-SB over existing methods grows even further.
Arithmetic Reasoning (Table 3): For Mistral-7B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on GSM8K by 1.90%, while achieving an impressive 234× reduction in communication cost. Additionally, Fed-SB ($r=200$) surpasses FFA-LoRA on GSM8K by 2.51%, with an approximately lower communication cost. Similarly, for Gemma-2 9B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on MATH by 0.86%, while reducing communication cost by 230×.
4.2 (Federated) Private Fine-Tuning
Details. We fine-tune BERT-base (Devlin, 2018) on SNLI (Bowman et al., 2015), a standard benchmark for natural language inference. Following Hu et al. (2021), we apply LoRA modules only to the self-attention layers. Our evaluation considers two DP settings: a centralized private setup and a federated private setup. To enforce DP guarantees during training, we use the Opacus library (Yousefpour et al., 2021) with the DP-SGD optimizer (Abadi et al., 2016). In the federated setting, training is conducted in a cross-silo setup with $3$ clients. We conduct experiments across a range of privacy budgets, varying $\epsilon$ from $1$ to $10$.
Results (Tables 4, 5). Fed-SB consistently outperforms all prior baselines in both accuracy and communication/parameter efficiency across all privacy budgets in both settings. Figures 4, 5, and 6 further illustrate this significant improvement.
Centralized Private (Table 4). We first evaluate our method in the centralized setting against existing approaches. Fed-SB consistently outperforms previous methods while utilizing only a fraction of the parameters across all $\epsilon$ values. For instance, at $\epsilon=3$, Fed-SB ($r=64$) surpasses centralized LoRA and centralized FFA-LoRA by 9.28% and 2.05%, respectively, while using approximately 12x and 6x fewer parameters.
Federated Private (Table 5). Fed-SB consistently outperforms all previous methods across all values of $\epsilon$, while significantly reducing communication costs. For instance, at $\epsilon=1$, Fed-SB ($r=64$) outperforms FedIT, FedEx-LoRA/FLoRA, and FFA-LoRA by 24.26%, 6.48%, and 2.72%, respectively, while reducing communication cost by approximately 12x, 35x, and 6x. FedIT performs significantly worse in the federated private setting compared to the federated non-private setting. We hypothesize that this is due to increased deviation in updates under DP constraints and added noise, leading to greater divergence from the ideal.
5 Conclusion
Existing LoRA-based federated FT methods either suffer from suboptimal updates or incur prohibitively high communication costs. We introduce Fed-SB, a federated adaptation of LoRA-SB that ensures exact aggregation while maintaining high communication efficiency. By training only a small $r\times r$ matrix and leveraging direct averaging, Fed-SB eliminates high-rank update costs and achieves communication efficiency independent of the number of clients. Fed-SB is particularly well-suited for private FT, as its linearity prevents noise amplification, and its reduced parameter count minimizes noise required for enforcing DP guarantees. It consistently outperforms existing methods across all models and tasks while reducing communication costs by up to 230x. These advantages establish Fed-SB as an efficient and scalable solution, setting a new Pareto frontier in (private) federated FT.
For future work, we plan to analyze the geometry of the $R$ matrices, as it is important for interpretability and may provide insights into the learning process. We also aim to study how the $R$ matrices change during training, which is key to understanding the fine-tuning process and characterizing its properties.
6 Limitations
While our approach is easily adaptable to other architectures, such as vision language models (VLMs) and vision transformers (VITs), we have not evaluated on such architectures. We have not yet extended our work to rank-heterogeneous settings, where clients operate with different ranks and computational budgets. Furthermore, we do not evaluate Fed-SB in scenarios with extreme data heterogeneity or significant class imbalances.
7 Acknowledgements
This work was supported by funding from Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) and an ADIA Lab fellowship.
Appendix A Proof of Lemma 1
Lemma.
Consider a model with $d$ learnable parameters trained using DP-SGD. The privacy parameter $\epsilon$ for $\delta$-approximate differential privacy, given $T$ training steps and a batch size of $q$, is expressed as: $\epsilon=O(q\sqrt{Td\log(1/\delta)})=O(\sqrt{d}).$ (13)
Proof.
The following result from Abadi et al. (2016) describes the relationship between noise variance, privacy parameters, number of optimization steps, batch size, and sample size in DP-SGD.
Theorem.
There exist constants $c_{1}$ and $c_{2}$ such that, given the sampling probability $q=L/N$ and the number of optimization steps $T$, for any $\epsilon<c_{1}q^{2}T$, DP-SGD is $(\epsilon,\delta)$-differentially private for any $\delta>0$ if the noise scale satisfies:
$\sigma\geq c_{2}\frac{q\sqrt{T\log(1/\delta)}}{\epsilon}.$ (14)
Each DP-SGD step introduces noise following $\mathcal{N}\left(0,\sigma^{2}C^{2}\mathbf{I}_{d}\right)$ and satisfies $(\alpha,\alpha/(2\sigma^{2}))$-RDP (Rényi DP) for the Gaussian mechanism. For a function with $\ell_{2}$-sensitivity $\Delta_{2}$, the Gaussian mechanism satisfies $(\alpha,\epsilon)$-RDP with:
$\epsilon(\alpha)=\frac{\alpha\Delta_{2}^{2}}{2\sigma_{\text{noise}}^{2}}.$ (15)
Since DP-SGD has $\Delta_{2}=C$ and $\sigma_{\text{noise}}=\sigma C$, applying privacy amplification due to sampling probability $q$ results in each step satisfying $(\alpha,\gamma)$-RDP, where, for small $q$:
$\gamma=O\left(\frac{q^{2}\alpha}{\sigma^{2}}\right).$ (16)
Using composition over $T$ steps, the total RDP privacy parameter becomes:
$\gamma_{\text{total}}=O\left(\frac{q^{2}T\alpha}{\sigma^{2}}\right).$ (17)
Converting this RDP bound back to $(\epsilon,\delta)$-DP and setting $\alpha$ proportional to $1/\sqrt{d}$, given that the $\ell_{2}$-norm of the gradient scales as $\sqrt{d}$, we obtain:
$\epsilon=O\left(\frac{q^{2}T\alpha}{\sigma^{2}}+\frac{\log(1/\delta)}{\alpha-1}\right).$ (18)
Substituting $\sigma\propto 1/\sqrt{d}$, we derive:
$\epsilon=O(q\sqrt{Td\log(1/\delta)})=O(\sqrt{d}).$ (19)
Appendix B Related Work
Parameter-Efficient Fine-Tuning (PEFT). LoRA (Hu et al., 2021) has become ubiquitous for fine-tuning LLMs (Zhang et al., 2024a) by modeling weight updates as product of low-rank matrices. Several variants have been proposed to improve efficiency, stability, and adaptability. QLoRA (Dettmers et al., 2024) enables efficient fine-tuning through quantization strategies, reducing memory usage while maintaining performance. AdaLoRA (Zhang et al., 2023) dynamically allocates a layer-specific rank budget by assigning importance scores to individual weight matrices. LoRA-XS (Bałazy et al., 2024) further reduces trainable parameters by inserting a trainable matrix between frozen LoRA matrices. VeRA (Kopiczko et al., 2024) enhances parameter efficiency by learning shared adapters across layers. DoRA (Liu et al., 2024) decomposes the pre-trained matrix into two parts—magnitude and direction—and applies LoRA modules only to the direction component. PiSSA (Meng et al., 2024) improves adaptation by initializing adapters using the singular value decomposition (SVD) of pre-trained weights. rsLoRA (Kalajdzievski, 2023) introduces a rank-scaling factor to stabilize learning. LoRA-SB (Ponkshe et al., 2024) provably approximates gradients optimally in low-rank spaces, achieving superior performance with significantly higher parameter efficiency.
Federated Fine-Tuning. Federated Learning (FL) consists of a centralized global model and multiple clients, each with its own local dataset and computational capacity. The global model is updated by aggregating client updates (Kairouz et al., 2021). FedBERT (Tian et al., 2022) focuses on federated pre-training, while other methods work on federated fine-tuning (Zhang et al., 2022; Kuang et al., 2024; Babakniya et al., 2023). Fed-IT (Zhang et al., 2024c) aggregates low-rank adapters across clients using standard federated averaging (McMahan et al., 2017) before updating the global model. To address inexact aggregation, FedEx-LoRA (Singhal et al., 2025) introduces an error matrix to correct residual errors, ensuring more precise updates. FLoRA (Wang et al., 2024) follows the same exact aggregation principle by stacking matrices and extends this approach to heterogeneous rank settings. FFA-LoRA (Sun et al., 2024) mitigates aggregation inexactness by freezing $\mathbf{A}$ and updating only the trainable low-rank adapter, averaging the latter to compute the global update. In some scenarios, clients require heterogeneous LoRA ranks due to varying computational budgets (Zhao et al., 2018; Li et al., 2019). Methods like HetLoRA (Cho et al., 2024) enable rank heterogeneity through self-pruning and sparsity-aware aggregation strategies, but incur significant overhead.
Differential Privacy (DP) and FL. A common limitation of standard FL frameworks is their susceptibility to privacy attacks, as clients publicly share model updates with a central server. To address this issue, DP is incorporated into FL methods to ensure the privacy of client updates. This work follows the approximate DP framework (Dwork, 2006; Dwork et al., 2014), which provides formal privacy guarantees for model updates. Privacy is enforced during training using the DP-SGD optimizer (Abadi et al., 2016), which applies gradient clipping and noise injection to protect individual contributions. Since DP is preserved under composition and post-processing (Dwork, 2006; Li et al., 2021), the final global model update also retains DP guarantees. Prior methods, such as Fed-IT and FedEx-LoRA, did not explicitly incorporate DP. This study extends these approaches to DP settings and benchmarks them alongside FFA-LoRA and the proposed method.
Appendix C Experiment Details
We conduct experiments on a single NVIDIA A6000 GPU (48 GB) and report the average results from three independent runs. All non-private models are trained using the AdamW optimizer (Loshchilov and Hutter, 2019). In line with Ponkshe et al. (2024), we initialize the adapter matrices using just $1/{1000}$ of the respective training dataset size.
Instruction Tuning. Table 6 presents the key hyperparameters and configurations for Mistral-7B, Gemma-2 9B, and Llama-3.2 3B. Our setup closely follows previous works (Hu et al., 2023; Ponkshe et al., 2024), ensuring consistency with established best practices. For the baseline experiments, we further set $\alpha=16$, consistent with prior literature (Singhal et al., 2025; Sun et al., 2024). We additionally perform a sweep over the learning rate for our experiments.
(Federated) Private Fine-Tuning. Table 7 outlines the key hyperparameters and configurations for BERT-base in both centralized private and federated private settings. We train our models using the Opacus library (Yousefpour et al., 2021) with the DP-SGD optimizer (Abadi et al., 2016). Following standard DP practices, we set the privacy parameter as $\delta=\frac{1}{|\text{trainset}|}$. To ensure adherence to best practices, we adopt hyperparameter choices from prior works (Singhal et al., 2025; Hu et al., 2021). For baseline experiments, we additionally set $\alpha=16$, aligning with previous literature (Singhal et al., 2025; Sun et al., 2024). We additionally perform a sweep over the learning rate and maximum gradient norm in DP-SGD for our experiments.
Mistral-7B Gemma-2 9B Llama-3.2 3B
Optimizer AdamW AdamW AdamW
Learning Rate $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $2\mathrm{e}{-4}$
LR Scheduler Cosine Cosine Linear
Warmup Ratio $0.02$ $0.02$ $0.02$
Batch Size $1$ $1$ $8$
Grad Acc. Steps $32$ $32$ $24$
Max. Seq. Len $512$ $512$ $256$
Dropout 00 00 00
# Clients $25$ $25$ $5$
Local Epochs $1$ $2$ $2$
Rounds $1$ $1$ $1$
Table 6: Hyperparameter settings for Mistral-7B, Gemma-2 9B, and Llama-3.2 3B.
BERT-base (centralized) BERT-base (federated)
Optimizer DP-SGD DP-SGD
Learning Rate $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$
LR Scheduler - -
Warmup Ratio 0 0
Batch Size $32$ $32$
Max. Phy. Batch Size $8$ $8$
Max. Seq. Len $128$ $128$
Dropout $0.05$ $0.05$
Max. Grad. Norm $0.1$ $0.1$
Epochs $3$
# Clients - $3$
Local Epochs - $6$
Rounds - $1$
Table 7: Hyperparameter settings for BERT-base in centralized private and federated private setups.
Appendix D Dataset Details
CommonSense170K is a large-scale dataset that brings together eight benchmarks designed to assess various aspects of commonsense reasoning (Hu et al., 2023). Below is an overview of its constituent datasets:
  1. 1.
    PIQA (Bisk et al., 2020) evaluates physical commonsense by asking models to determine the most reasonable action in a given scenario.
  2. 2.
    ARC Easy (ARC-e) (Clark et al., 2018) consists of elementary-level science questions, serving as a fundamental test of a model’s reasoning abilities.
  3. 3.
    OBQA (Mihaylov et al., 2018) presents knowledge-intensive, open-book multiple-choice questions that require multi-step reasoning and retrieval.
  4. 4.
    HellaSwag (Zellers et al., 2019) tests contextual reasoning by asking models to predict the most plausible continuation of a passage from a set of candidates.
  5. 5.
    SIQA (Sap et al., 2019) examines social intelligence, requiring models to predict human actions and their social consequences.
  6. 6.
    ARC Challenge (ARC-c) (Clark et al., 2018) includes difficult multiple-choice science questions that demand deeper logical inference beyond statistical co-occurrence.
  7. 7.
    BoolQ (Clark et al., 2019) consists of naturally occurring yes/no questions, requiring models to infer relevant information from provided contexts.
  8. 8.
    WinoGrande (Sakaguchi et al., 2021) assesses commonsense knowledge through binary-choice sentence completion tasks that require resolving ambiguities.
The MetaMathQA dataset (Yu et al., 2024) constructs mathematical questions by reformulating them from different viewpoints while preserving their original knowledge content. We assess its performance using two well-established benchmarks: (1) GSM8K (Cobbe et al., 2021), a collection of grade-school-level math problems requiring step-by-step reasoning to reach a solution, and (2) MATH (Hendrycks et al., 2021), which consists of high-difficulty, competition-style problems designed to test advanced mathematical skills.
Stanford Natural Language Inference (SNLI) is a widely used benchmark for assessing textual entailment models in natural language understanding. It contains approximately 570,000 sentence pairs, each categorized into one of three classes: entailment, contradiction, or neutral, requiring models to infer the relationship between a given premise and hypothesis.
Appendix E Additional Plots
((a)) Mistral-7B (GSM8K)
((b)) Gemma-2 9B (MATH)
((c)) Llama-3.2 3B (Commonsense)
Figure 3: Performance vs. number of communicated parameters (in log scale) for various methods in federated fine-tuning across multiple models on arithmetic and commonsense reasoning tasks.
((a)) Centralized Private
((b)) Federated Private
Figure 4: Performance comparison of various methods in centralized (Cent.) private and federated private fine-tuning (BERT-base) on SNLI across varying values of $\epsilon$.
((a)) $\epsilon=1$
((b)) $\epsilon=3$
((c)) $\epsilon=5$
((d)) $\epsilon=7.5$
((e)) $\epsilon=10$
Figure 5: Performance vs. number of trainable parameters (in log scale) for various methods in centralized private fine-tuning (BERT-base) across different privacy budgets ($\epsilon$).
((a)) $\epsilon=1$
((b)) $\epsilon=3$
((c)) $\epsilon=5$
((d)) $\epsilon=7.5$
((e)) $\epsilon=10$
Figure 6: Performance vs. number of communicated parameters (in log scale) for various methods in federated private fine-tuning (BERT-base) across different privacy budgets ($\epsilon$).